# Learning Variable Compliance Control From a Few Demonstrations for Bimanual Robot with Haptic Feedback Teleoperation System

### v1

@@ Abstract
Automating dexterous, contact-rich manipulation tasks using rigid robots is a significant challenge in robotics. Rigid robots, defined by their actuation through position commands, face issues of excessive contact forces due to their inability to adapt to contact with the environment, potentially causing damage. While compliance control schemes have been introduced to mitigate these issues by controlling forces via external sensors, they are hampered by the need for fine-tuning task-specific controller parameters. Learning from Demonstrations (LfD) offers an intuitive alternative, allowing robots to learn manipulations through observed actions. In this work, we introduce a novel system to enhance the teaching of dexterous, contact-rich manipulations to rigid robots. Our system is twofold: firstly, it incorporates a teleoperation interface utilizing Virtual Reality (VR) controllers, designed to provide an intuitive and cost-effective method for task demonstration with haptic feedback. Secondly, we present Comp-ACT (Compliance Control via Action Chunking with Transformers), a method that leverages the demonstrations to learn variable compliance control from a few demonstrations. Our methods have been validated across various complex contact-rich manipulation tasks using single-arm and bimanual robot setups in simulated and real-world environments, demonstrating the effectiveness of our system in teaching robots dexterous manipulations with enhanced adaptability and safety.
@@ I Introduction
Learning from Demonstrations (LfD) has emerged as an intuitive and promising solution to teach robots dexterous manipulations [3, 4]. Instead of manually programming a sequence of actions or engineering a reward function for the robot to learn from, in LfD, the operator can show the robot how exactly to perform the task. However, the challenges of effectively using LfD are sample efficiency-the need of many demonstrations for each task-, selecting optimal stiffness parameters for the compliance controller, and the teaching interface. We propose a robot demonstration system for contact-rich manipulation tasks to address these challenges.
@@
Secondly, inspired by [11], we propose a method to learn variable Compliance Control via Action Chunking with Transformers (Comp-ACT). Comp-ACT consists of gathering demonstrations, including the Cartesian trajectory of the robots’ end-effector, the measured F/T on each robot, the compliance control parameters during teleoperation (i.e., stiffness), and camera images from multiple points of view. Following that, we train a Comp-ACT policy to predict a chunk of future actions given the current state of the robot. The main differences with [11] are that we learn a policy using the Cartesian task space instead of joint space; the external F/T is also used to condition the future action; and the stiffness parameters of the compliance controller are simultaneously learned. In summary, we make the following contributions;
@@ II-B Teleoperation Systems
Most of the previously proposed VR-based teleoperation systems implemented a virtual world where the human operator is aligned with the simulated robot. For simplicity, this work uses only the VR controllers’ tracking capability. Our approach allows for easier adoption of our system with different robots. Additionally, we use haptic feedback as it has been shown to improve robotic arm control [28]. Our haptic feedback interface uses the VR controller’s vibration mechanism, similar to [29], where they use the phone vibrations.
@@ III-A Teleoperation System
Our teleoperation system works as follows. First, we use PyOpenVR [30], a Python wrapper for ValveSoftware’s OpenVR library, to obtain the tracked pose of the controllers at 90Hz. We are interested in the relative translation and orientation of the controller from a defined starting pose. The relative motion of the controller is then mapped to the pose of the robot’s end effector (EE) and defined as the target pose for the controller to follow. The operator sets the starting pose of the controller used to compute the relative motion through the menu button of the controller, which also enables/disables the teleoperation. This control mechanism allows the user to operate the robot easily regardless of the workspace size difference due to the different lengths of the human arm and the robot arm. Additionally, the trigger on the controller is used to signal the desired gripper pose, which ranges from zero to a hundred percent of the maximum opening width of the gripper. The grip button switches between different degrees of compliance, i.e., the compliance controller stiffness. The haptic feedback to the user is achieved by mapping the contact force measure at the robot’s wrist F/T sensor to the level of intensity of the controller’s vibration. The button bindings are shown in Figure 3.
@@ IV-A Transformer-based Behavioral Cloning
The Transformer-based policy has shown great promise in processing long-horizon multimodal action data in an autoregressive manner. We built Comp-ACT upon Action Chunking with Transformers [11] to learn Cartesian EE trajectory and robots’ stiffness from demonstrations. ACT predicts a sequence of future actions instead of a single action like in standard Behavioral Cloning (BC), mitigating compounding errors. Following ACT, we learn the policy as the decoder of a conditional variational autoencoder (CVAE). It predicts a sequence of action conditioned on the current observation, including RGB images, F/T sensor readings, and proprioception. Figure 4 illustrates the architecture of the CVAE. The predicted Cartesian EE pose and stiffness are then fed into the compliance controller described in Section III-B. Each robot’s action includes three components: a 6-dimensional Cartesian pose (with position and orientation in 3D space, the latter in axis-angle form), a 1-dimensional gripper action (non-binary desired opening width), and a 12-dimensional vector to represent the stiffness parameters. Consequently, the bimanual robot’s action space is 38-dimensional. We use the Cholesky representation to encode the stiffness matrix for translation and rotation, similar to [33]. The Cholesky representation is the flattened form of the upper triangular matrix $\hat{K}$ from the Cholesky decomposition of the stiffness matrix $K$, where $K=\hat{K}^{T}\hat{K}$.
IV-B Data Collection
V Experiments
V-A Experimental Setup
We used two UR5e robot arms with built-in force-torque sensors in simulation and real-world experiments, with bimanual and single-arm tasks. Two cameras, Realsense SR305, are attached to the wrist of each arm. Another static camera with a wider field of view is in front of the task space. Different combinations of these cameras were used as observations on each task and described in Section V-B. Figure 1 shows our experimental setup.
For simulation, we use the physics engine MuJoCo [34]. A virtual twin of our robotic setup was created as a new robosuite [35] environment, as shown in Figure 8.
The robotic system runs on a single machine with a CPU Intel i7-12700F and a GPU NVIDIA GeForce RTX 3070. The simulation and the policy training are performed on a different machine with a GPU NVIDIA GeForce RTX 4090.
TASK DESCRIPTION
Figure 9: Bimanual Peg-in-Hole (Cylinder): A 3D-printed peg and matching hole with tolerance of $2$ mm were used for this task. The relative position and orientation of the two arms need to be accurately aligned to complete the round-shaped peg insertion. During demonstrations, the demonstrator operates the left arm in a medium stiffness mode for free motion and switches to high stiffness to make the robot resist motion when the other arm pushes the peg into the hole. On the contrary, the right arm switches to low stiffness to allow contact for fine adjustment of the peg’s position. 20 demonstrations were collected for this task.
Figure 10: Bimanual Peg-in-Hole (Cuboid): A 3D-printed peg and matching hole with tolerance of $2$ mm were used for this task. Besides matching the peg and hole’s surface position and orientation, the cuboid insertion also requires aligning each corner by rotating around the direction perpendicular to the surface. The same stiffness switching strategy described in Figure 10 was used during demonstrations. 30 demonstrations were collected for this task.
V-B Tasks
@@ VI-A Simulation Task
We evaluated our proposed method Comp-ACT compared to ACT [11] on simulation with the bimanual wiping task described in Figure 8. The same demonstrations were used to compare both methods fairly. For ACT, the action was defined as the joint positions of each robot, and we confirmed that replaying the demonstrations using joint-position control achieves similar contact forces. The learned policies were tested 10 times each. Both policies learned and solved the task correctly, i.e., the hammer is picked and lifted, the mark on the table wiped, and then the hammer is returned to its original place. However, as shown in Figure 11, when we examined the contact force applied to the table, it is clear that when using joint-position control with the ACT policy, the robot would apply force to the table of more than 5 times higher than when using our proposed Comp-ACT method. In the case of ACT, minor errors in the trajectory predicted by the policy would result in significant forces being applied to the table due to the joint-position control scheme. In contrast, with our compliance control scheme, minor errors in the trajectory would not result in significant forces due to the low stiffness used while wiping.
@@
Figure 11: Contact force comparison for bimanual wiping task in simulation. Only the contact force measured by the wiping robot is considered here. The plot shows the average contact force (continuous line) of 10 trials for each learned policy and the demonstrations, as well as the standard deviation (shaded colored area).
Additionally, we compared our proposed method that includes the F/T data as of the observation with a version that did not include it, called Comp-ACT w/o F/T. Figure 11 shows that including the F/T resulted in more consistently applying less contact force during the wiping action. Nevertheless, even without including the F/T data, the contact force was similar to or even lower than that achieved by the demonstrator.
@@
Figure 12: Comparison of drawings. Sampled images (Top) and average images (Bottom) from demonstrations (Left), Comp-ACT (Middle), and Comp-ACT without F/T data (Right).
VI-B Real-World Tasks
Picking & Insertion
This task is challenging because a successful insertion depends greatly on the orientation of the peg after picking. Comp-ACT policies, both with and without F/T, consistently succeeded in picking the peg. The success rates for the insertion phase were 70% with F/T and 35% without, indicating the effectiveness of F/T data in enhancing peg-in-hole task performance. Despite the challenges presented by the variability in the peg’s starting position, the different orientations from the picking phase during demonstrations, and the movable wooden box, achieving a 70% success rate in insertion with only 30 demonstrations is remarkable.
Wiping
Drawing
In this case, success was defined subjectively by the resemblance of the drawing to an emoji smile. Figure 12 shows the drawings by demonstrators and trained policies (with and without F/T). All trials successfully produced emoji smiles with accurately placed eyes and mouths, using appropriate force for clear, continuous lines. Qualitative analysis of similarity to the demonstrated smiles using the Structural Similarity Index (SSIM) [36] showed negligible differences in performance between Comp-ACT with an average similarity score of 84.2% and Comp-ACT without F/T at 84.3%. Similarly to the wiping task, policies without F/T data exhibited longer motions.
Peg-in-Hole tasks
The two bimanual peg-in-hole insertion tasks are the most challenging of the ones considered in this paper. While the robot’s initial position remained the same in each demonstration, the pose where the robots joined to perform the insertion changed every time, as the demonstrators could not reproduce the same motions perfectly. Nevertheless, we observed high success rates for both tasks with our proposed method: 100% and 70% for the cylindrical peg insertion and the cuboid peg insertion, respectively. The success rate drops significantly when the F/T data is not included in the Comp-ACT policy: 40% and 50% for cylindrical and cuboid peg insertion, respectively. These results further corroborate the importance of F/T data for peg insertion tasks. In both tasks, the failure cases observed mainly consisted of misalignment of the parts and subsequently applying excessive contact force or releasing the peg too early before the insertion.
@@ VIII Conclusion
Additionally, we showed why the baseline ACT method [11], with its joint position-based control, would not be suitable for rigid robots as errors in the predicted trajectory could result in undesired large contact forces that may damage the manipulated objects or the robot. On the contrary, Comp-ACT reduces the risk of excessive forces by learning the stiffness parameters of the underlying compliance controller, enabling safe task execution with variable compliance in Cartesian space.

### v2

@@ Abstract
Automating dexterous, contact-rich manipulation tasks using rigid robots is a significant challenge in robotics. Rigid robots, defined by their actuation through position commands, face issues of excessive contact forces due to their inability to adapt to contact with the environment, potentially causing damage. While compliance control schemes have been introduced to mitigate these issues by controlling forces via external sensors, they are hampered by the need for fine-tuning task-specific controller parameters. Learning from Demonstrations (LfD) offers an intuitive alternative, allowing robots to learn manipulations through observed actions. In this work, we introduce a novel system to enhance the teaching of dexterous, contact-rich manipulations to rigid robots. Our system is twofold: firstly, it incorporates a teleoperation interface utilizing Virtual Reality (VR) controllers, designed to provide an intuitive and cost-effective method for task demonstration with haptic feedback. Secondly, we present Comp-ACT (Compliance Control via Action Chunking with Transformers), a method that leverages the demonstrations to learn variable compliance control from a few demonstrations. Our methods have been validated across various complex contact-rich manipulation tasks using single-arm and bimanual robot setups in simulated and real-world environments, demonstrating the effectiveness of our system in teaching robots dexterous manipulations with enhanced adaptability and safety. Code available at https://github.com/omron-sinicx/CompACT.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Learning from Demonstrations (LfD) has emerged as an intuitive and promising solution to teach robots dexterous manipulations  [3]. Instead of manually programming a sequence of actions or engineering a reward function for the robot to learn from, in LfD, the operator can show the robot how exactly to perform the task. However, the challenges of effectively using LfD are sample efficiency-the need for many demonstrations for each task-, selecting optimal stiffness parameters for the compliance controller, and the teaching interface. We propose a robot demonstration system for contact-rich manipulation tasks to address these challenges.
@@
Secondly, inspired by [10], we propose a method to learn variable Compliance Control via Action Chunking with Transformers (Comp-ACT). It starts by gathering demonstrations, including the Cartesian trajectory of the robots’ end-effector, the measured F/T on each robot, the compliance control parameters during teleoperation (i.e., stiffness), and camera images from multiple points of view. Following that, we train a Comp-ACT policy to predict a chunk of future actions given the current state of the robot. The main differences with [10] are that we learn a policy using the Cartesian task space instead of joint space; the external F/T is also used to condition the future action; and the stiffness parameters of the compliance controller are simultaneously learned. In summary, we make the following contributions;
@@ II-B Teleoperation Systems
Most of the previously proposed VR-based teleoperation systems implemented a virtual world where the human operator is aligned with the simulated robot. For simplicity, this work uses only the VR controllers’ tracking capability. Authors in [28] proposed a tele-impedance framework using a custom-designed interface similar to VR controller, validated on a torque-controlled robot. Our teleoperation interface, on the other hand, adopts a commercially available VR controller combined with a compliance controller, which is required for adopting the system for rigid robots. Additionally, we use haptic feedback as it has been shown to improve robotic arm control [29]. Our haptic feedback interface uses the VR controller’s vibration mechanism, similar to [30], where they use the phone vibrations.
@@ III-A Teleoperation System
Our teleoperation system works as follows. First, we use PyOpenVR [31], a Python wrapper for ValveSoftware’s OpenVR library, to obtain the tracked pose of the controllers at 90Hz. We are interested in the relative translation and orientation of the controller from a defined starting pose. The relative motion of the controller is then mapped to the pose of the robot’s end-effector (EE) and defined as the target pose for the controller to follow. The operator sets the starting pose of the controller used to compute the relative motion through the menu button of the controller, which also enables/disables the teleoperation. This control mechanism allows the user to operate the robot easily regardless of the workspace size difference due to the different lengths of the human arm and the robot arm. Additionally, the trigger on the controller is used to signal the desired gripper pose, which ranges from zero to a hundred percent of the maximum opening width of the gripper. The grip button switches between different degrees of compliance, i.e., the compliance controller stiffness. The haptic feedback to the user is achieved by mapping the contact force measure at the robot’s wrist F/T sensor to the level of intensity of the controller’s vibration. The button bindings are shown in Figure 3.
@@ IV-A Transformer-based Behavioral Cloning
The Transformer-based policy has shown great promise in processing long-horizon multimodal action data in an autoregressive manner. We built Comp-ACT upon Action Chunking with Transformers [10] to learn Cartesian EE trajectory and robots’ stiffness from demonstrations. ACT predicts a sequence of future actions instead of a single action like in standard Behavioral Cloning (BC), mitigating compounding errors. Following ACT, we learn the policy as the decoder of a conditional variational autoencoder (CVAE). It predicts a sequence of action conditioned on the current observation, including RGB images, F/T sensor readings, and proprioception. Figure 4 illustrates the architecture of the CVAE. The predicted Cartesian EE pose and stiffness are then fed into the compliance controller described in Section III-B. Each robot’s action includes three components: a 6D Cartesian pose (with position and orientation in 3D space, the latter in axis-angle form), a 1D gripper action (non-binary desired opening width), and a 12D vector to represent the stiffness parameters. Consequently, the bimanual robot’s action space is 38-dimensional. We use the Cholesky representation to encode the stiffness matrix for translation and rotation for better extensibility to full stiffness matrix, similar to [34]. The Cholesky representation is the flattened form of the upper triangular matrix $\hat{K}$ from the Cholesky decomposition of the stiffness matrix $K$, where $K=\hat{K}^{T}\hat{K}$.
@@
IV-B Data Collection
V Experiments
V-A Experimental Setup
We used two UR5e robot arms with built-in force-torque sensors in simulation and real-world experiments, with bimanual and single-arm tasks. Two cameras, Realsense SR305, are attached to the wrist of each arm. Another static camera with a wider field of view is in front of the task space. Different combinations of these cameras were used as observations on each task and described in Section V-B. Figure 1 shows our experimental setup. For simulation, we use the physics engine MuJoCo [35]. A virtual twin of our robotic setup was created as a new robosuite [36] environment, as shown in Figure 8. The robotic system runs on a single machine with a CPU Intel i7-12700F and a GPU NVIDIA GeForce RTX 3070. The simulation and the policy training are performed on a different machine with a GPU NVIDIA GeForce RTX 4090.
TASK DESCRIPTION
Figure 9: Bimanual Insertion (Cylinder): A 3D-printed peg and matching hole with tolerance of $2$ mm. The relative position and orientation of the two arms need to be accurately aligned to complete the round-shaped peg insertion. During demonstrations, both arms start with medium stiffness mode during contact-free motion. The left arm switches to high stiffness to resist motion during the insertion. On the contrary, the right arm switches to low stiffness to slow the robot and allow force to guide the insertion. 20 demonstrations were collected for this task.
Figure 10: Bimanual Insertion (Cuboid): A 3D-printed peg and matching hole with tolerance of $2$ mm. Besides matching the peg and hole’s surface position and orientation, the cuboid insertion also requires aligning each corner by rotating around the direction perpendicular to the surface. The same stiffness switching strategy described in Figure 10 was used during demonstrations. 30 demonstrations were collected for this task.
V-B Tasks
@@ VI-A Simulation Task
We evaluated our proposed method Comp-ACT compared to ACT [10] on simulation with the bimanual wiping task described in Figure 8. The effect of adding F/T data as observation was also investigated. The same demonstrations were used to compare both methods fairly. For ACT, the action was defined as the joint positions of each robot, and we confirmed that replaying the demonstrations using joint-position control achieves similar contact forces. The learned policies were tested 10 times each.
VI-A1 Comp-ACT vs ACT
Both policies learned and solved the task correctly, i.e., the hammer is picked and lifted, the mark on the table wiped, and then the hammer is returned to its original place. However, as shown in Figure 11, when we examined the contact force applied to the table, it is clear that when using joint-position control with the ACT policy, the robot would apply force to the table of more than 5 times higher than when using our proposed Comp-ACT method. In the case of ACT, minor errors in the trajectory predicted by the policy would result in significant forces being applied to the table due to the joint-position control scheme. In contrast, with our compliance control scheme, minor errors in the trajectory would not result in significant forces due to the low stiffness used while wiping.
@@ VI-A1 Comp-ACT vs ACT
Figure 11: Force comparison for the simulated bimanual wiping task. Only the contact force measured by the wiping robot is considered. The average contact force (continuous line) and standard deviation (shaded colored area) of 20 trials for each policy and the demonstrations are shown.
VI-A2 F/T as observation vs Not
Adding F/T data to the observation of ACT policy (ACT + F/T) helped suppress the applied force compared to ACT without F/T observation (ACT), yet both applied excessive force to the table. However, there was no significant difference in the force plot between Comp-ACT and Comp-ACT without F/T. We argue that it is because simply following the predicted reference Cartesian trajectory with a specified stiffness should already achieve relatively lower contact force similar to the demonstration, thus the F/T data did not have much information to the policy for this wiping task.
@@ VI-A2 F/T as observation vs Not
Figure 12: Comparison of drawings. Averaged images from demonstrations (Left), Comp-ACT (Middle), and Comp-ACT without F/T data (Right).
@@
VI-B Real-World Tasks
Picking & Insertion
This task is challenging because a successful insertion depends greatly on the orientation of the peg after picking. Comp-ACT policies, both with and without F/T, consistently succeeded in picking the peg. The success rates for the insertion phase were 70% with F/T and 35% without, indicating the effectiveness of F/T data in enhancing peg-in-hole task performance.
Wiping
Drawing
In this case, success was defined subjectively by the resemblance of the drawing to an emoji smile. Figure 12 shows the drawings by demonstrators and trained policies (with and without F/T). All trials successfully produced emoji smiles with accurately placed eyes and mouths, using appropriate force for clear, continuous lines. Quantitative analysis of similarity to the demonstrated smiles using the Structural Similarity Index (SSIM) [37] showed negligible differences in performance between Comp-ACT with an average similarity score of 84.2% and Comp-ACT without F/T at 84.3%. Similarly to the wiping task, policies without F/T data exhibited longer motions. Furthermore, the policy’s predicted stiffness action versus the normal force applied during the drawing task is visualized in Figure 13. The policy successfully chooses lower stiffness while in contact with the whiteboard, which realizes avoiding excessive force.
Peg-in-Hole tasks
The two bimanual peg-in-hole insertion tasks are the most challenging of the ones considered in this paper. While the robot’s initial position remained the same in each demonstration, the pose where the robots joined to perform the insertion changed every time, as the demonstrators could not reproduce the same motions perfectly. Nevertheless, we observed high success rates for both tasks with our proposed method: 100% and 70% for the cylindrical peg insertion and the cuboid peg insertion, respectively. The success rate drops significantly when the F/T data is not included in the Comp-ACT policy: 40% and 50% for cylindrical and cuboid peg insertion, respectively. These results further corroborate the importance of F/T data for peg insertion tasks. In both tasks, the failure cases observed mainly consisted of misalignment of the parts and subsequently applying excessive contact force or releasing the peg too early before the insertion. Videos of the policy rollouts including the failure cases can be found in our project website.
Figure 13: Drawing task: contact force versus predicted stiffness action during policy evaluation.
@@ VIII Conclusion
Additionally, we showed that the baseline ACT method [10], with its joint position-based control, would not be suitable for rigid robots as errors in the predicted trajectory could result in undesired large contact forces that may damage the manipulated objects or the robot. On the contrary, Comp-ACT reduces the risk of excessive forces by learning the stiffness parameters of the underlying compliance controller, enabling safe task execution with variable compliance in Cartesian space.